### v3

@@ A.3 Gradient calculations of $V_{th,l}^{1}$ and $S_{q}^{l}$
We adopt the same gradient calculations of the quantization step size of [13].
@@
where the minimum quantization integer $q_{min}=0$ and the maximum integer $q_{max}=2^{B_{s,l}}-1$. While, $g_{s,scale}$, equal to $\frac{1}{\sqrt{\sum_{j}1\cdot q_{max}}}$ and meant to avoid gradient explosion [13], is the gradient scaling factor. $L_{task}$ represents the task loss signal.
@@ A.5.1 CIFAR
In the CIFAR experiments, two models are adopted ResNet20 and Spikformer. For ResNet20, we adhere to the default settings of [18] and [19]. The training epochs are set to 200, batch size to 128, and learning rate to 0.1. Besides, the penalty coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 4e-2, 4e-2, and 1e-2, respectively. The initial W/S/T are set to 4/4/2. These settings apply to both comparative and ablation studies. For Spikformer, we adopt the same default settings as [56] from their official codes, including every training configuration. The penalty coefficients and the initial W/S/T are the same as our ResNet setting.
@@ A.5.3 CIFAR10-DVS
In the CIFAR10-DVS experiments, we only use the default DVS Spikformer defined by the official codes of [56]. To better observe the temporal dimension’s influence on the model accuracy of the event dataset. We set $T_{l}$ back to the unlearnable constant and the corresponding temporal squeezing is also deactivated. The corresponding penalty coefficients $\lambda_{1}$ and $\lambda_{3}$ are all set to 4e-2. The target weight bit widths are all set to 1, the target spike bit widths are 2 or 4. Under this setting, SNNs can fully exploit the model-level sequential processing ability. Thus, we can accurately observe that our methods can mitigate "the temporal constraint" steadily. In Section A.7, we also list the results of applying the learnable temporal lengths and temporal squeezing.
@@ A.5.4 DVS-GESTURE
In the DVS-GESTURE experiments, we also use the default DVS Spikformer defined by the official codes of [56]. Most experimental settings are the same as the CIFAR10-DVS experiments and the official codes [56]. Following is some exceptions: the training epoch is set to 192, and when the target time length is 1, 2, or x (x>2), each event data is divided into 4, 10, or 16 temporal slices, respectively. Only the first 1, 2, or x temporal slices are extracted to training and inference.
A.5.5 SHD
We follow the exactly same hyperparameters and model architectures of DCLS [20] without changing a thing. The corresponding penalty coefficient $\lambda_{1}$ is set to 4e-2. The initial W is set to 6 and target Ws are set to 3 and 4 in Table 8. The initial and target Ss are all set to 1. The initial and target Ts are all set to 140.
A.5.6 Additional statements about the quantization baselines
The closest baseline-work to ours is [39]. However, we find that their work is not open-sourced and lacks a lot of implementation details. Therefore, we try to reproduce their method based on their paper and our practice as we call "U-quant." in the main text. Nevertheless, we also find many of their statistics, such as S-ACE, may be roughly calculated so we have to re-calculate every statistics we show in the main text with more accurate means.
@@ A.7 Additional experimental results on ImageNet and CIFAR10-DVS
Firstly, as described in Section A.5, we use the initial T=1 in the main text. This may raise the concern about whether the temporal squeezing cannot function well on such up-scaled dataset. As shown in Table A.1, with the temporal squeezing, the model can achieve the same model accuracy while using fewer S-ACEs, handling the change of temporal length from 2 to 1 during training well. As for the reason why the model initialized as W/S/T=4/4/2 can perform better on accuracy very slightly. One explanation is that the initial temporal length does not affect the model performance while the target temporal length (or the actual temporal length) does, which is in line with the prior finding of [39] that the temporal length rarely affects the performance of multi-bit SNNs on static datasets. Therefore, for the saving of training time, we choose to set the initial temporal length as 1 in the experiments of the main text since the longer initial temporal length causes longer forward and backward inference time. Additionally, from another perspective, this could further prove that our multi-bit SNN has advanced its spatial processing ability to the limit, one time-step would suffice to fulfill the requirement of a high-accuracy inference.
In addition, we also naively apply our methods to Spikformer-8-512 on ImageNet with the same model hyperparameters as [39] and [56]. Their models achieve 70.87% [39] and 73.38% [56] with W/S/T = 4/2/1 and 16/1/4, respectively. While, our Spikformer achieves a higher 74.87% top-1 accuracy with lower bit budgets that W/S/T = 2.64/2.21/1.0.
@@
Secondly, as stated in Section A.5, we use the conventional sequential-processing pipeline to address the event dataset in the main text, i.e., the temporal squeezing and learnable $T$ are canceled. In this way, we can accurately control the temporal length and fully utilize the model-level temporal processing of SNN. This may raise the concern that the temporal squeezing could severely impair the temporal processing ability. As shown in Table A.2, we find the temporal squeezing can still outperform the quantization baseline [39] and maintain good performance under low time-steps, although the overall model performance, compared to the case without temporal squeezing, is decreased. One possible explanation is that although the model-level temporal processing ability may be diminished, the layer-wise temporal processing ability can still be maintained when using temporal squeezing. And, temporal squeezing also meets the temporal information dynamics that as the training goes on, more valuable and decisive information aggregates into the early time-steps [26]. As listed in Table A.2, the best results of temporal squeezing show up under the low time-step settings. Temporal squeezing merely accelerates the aggregation process. That is why temporal squeezing would not severely undermine the sequential-processing ability and perform well even on the sequential dataset.
@@
Origin[39] Quant. Spf.[39] Our Spf. Our Spf. w/ Learnable $T_{l}$ W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 80.7 1/1/16 79.8 1.11/2.00/6.0 80.9 1.14/1.97/6.00 78.7 1/2/8 79.3 1.28/2.00/4.0 80.0 1.14/1.97/4.00 79.3 1/4/4 63.1 1.33/4.00/3.0 80.1 1.17/4.08/2.05 79.6 1/8/2 43.0 1.31/4.03/2.0 78.7 1.09/1.95/1.95 80.0 1/16/1 35.8 1.31/4.03/1.0 77.4 1.11/1.49/1.00 77.3
@@
A.8 Maintaining the asynchronous nature of SNNs
@@
Exploiting the spike perceptrons. Temporal squeezing itself is consistent with neurodynamics as it is essentially composed of a signal accumulation [17] and a division operation. A similar accumulation mechanism can also be found in the current accumulation stage of spiking neuron, such as LIF and IF. Particularly, SpikeConverter neuron [32] separates this accumulation from the spiking neuron modeling and uses a particular computing pipeline to handle this accumulation. Their accumulation is identical to ours, which is proven feasible in asynchronous computing. In addition, such accumulation has also been realized and implemented by neuromorphic circuits to process the DVS signals and generate the dynamic datasets [2, 31].
As such, we can simply use the same signal accumulation diagrams, i.e. the spike perceptrons, to realize temporal squeezing as what has been used in DVS cameras [28] and previous spiking neurons [32]. In other words, such different computing paradigm as shown in Figure A.1 can be implemented under the existing neuromorphic computing and will not cause any loss of the asynchronous and clock-free nature.
Using the two-stage pipeline or keeping single time-step. Although we have theoretically elaborated on the harmlessness of temporal squeezing to the asynchronous nature in the above, one may argue that the entire network’s spike flow could be interrupted by the squeezing stage, leading to longer inference latency. This concern has already discussed and resolved by [32], using the two-stage pipeline. The first stage is spike accumulation, i.e., the temporal squeezing, the second stage is spike emission. These two separate stages can overlapped inter-layer. Specifically, assuming spike accumulation and emission both require 4 hardware clock cycles for computing. After the first cycle of the spike emission of the previous layer, the first computing cycle of the spike accumulation of the next layer can be started instead of waiting for the whole 4 cycles of the spike emission of the previous layer being finished. Thus, 3 cycles are overlapped, and the actual latency of the two-stage spiking pipeline is only 5 cycles, only one more cycle that the ordinary spiking pipeline. Therefore, using temporal squeezing only introduces very limited latency increase.
A.9 Overall algorithm of the proposed bit allocation

### v4

@@ 7 Conclusion
Finally, we provide some discussions and analyses concerning the potential hardware-level issues, including how to maintain the neuromorphic computing nature, hardware considered energy merit analysis, reusable neural integration unit design, and energy consumption analysis of individual synaptic operation.
@@ A.1 Guideline
    Section A.8: the potential exploration on spiking audio task.
  8. 8.
@@ A.3 Gradient calculations of $V_{th,l}^{1}$ and $S_{q}^{l}$
We adopt the same gradient calculations of the quantization step size of [esser2020learned].
@@
where the minimum quantization integer $q_{min}=0$ and the maximum integer $q_{max}=2^{B_{s,l}}-1$. While, $g_{s,scale}$, equal to $\frac{1}{\sqrt{\sum_{j}1\cdot q_{max}}}$ and meant to avoid gradient explosion [esser2020learned], is the gradient scaling factor. $L_{task}$ represents the task loss signal.
@@ A.5.1 CIFAR
In the CIFAR experiments, two models are adopted ResNet20 and Spikformer. For ResNet20, we adhere to the default settings of [guo2024ternary] and [guo2022real]. The training epochs are set to 200, batch size to 128, and learning rate to 0.1. Besides, the penalty coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 4e-2, 4e-2, and 1e-2, respectively. The initial W/S/T are set to 4/4/2. These settings apply to both comparative and ablation studies. For Spikformer, we adopt the same default settings as [zhou2022spikformer] from their official codes, including every training configuration. The penalty coefficients and the initial W/S/T are the same as our ResNet setting.
@@ A.5.3 CIFAR10-DVS
In the CIFAR10-DVS experiments, we only use the default DVS Spikformer defined by the official codes of [zhou2022spikformer]. To better observe the temporal dimension’s influence on the model accuracy of the event dataset. We set $T_{l}$ back to the unlearnable constant and the corresponding temporal squeezing is also deactivated. The corresponding penalty coefficients $\lambda_{1}$ and $\lambda_{3}$ are all set to 4e-2. The target weight bit widths are all set to 1, the target spike bit widths are 2 or 4. Under this setting, SNNs can fully exploit the model-level sequential processing ability. Thus, we can accurately observe that our methods can mitigate "the temporal constraint" steadily. In Section A.7, we also list the results of applying the learnable temporal lengths and temporal squeezing.
@@ A.5.4 DVS-GESTURE
In the DVS-GESTURE experiments, we also use the default DVS Spikformer defined by the official codes of [zhou2022spikformer]. Most experimental settings are the same as the CIFAR10-DVS experiments and the official codes [zhou2022spikformer]. Following is some exceptions: the training epoch is set to 192, and when the target time length is 1, 2, or x (x>2), each event data is divided into 4, 10, or 16 temporal slices, respectively. Only the first 1, 2, or x temporal slices are extracted to training and inference.
A.5.5 Additional statements about the quantization baselines
The closest baseline-work to ours is [shen2024conventional]. However, we find that their work is not open-sourced and lacks a lot of implementation details. Therefore, we try to reproduce their method based on their paper and our practice as we call "U-quant." in the main text. Nevertheless, we also find many of their statistics, such as S-ACE, may be roughly calculated so we have to re-calculate every statistics we show in the main text with more accurate means.
@@ A.7 Additional experimental results on ImageNet and CIFAR10-DVS
Firstly, as described in Section A.5, we use the initial T=1 in the main text. This may raise the concern about whether the temporal squeezing cannot function well on such up-scaled dataset. As shown in Table A.1, with the temporal squeezing, the model can achieve the same model accuracy while using fewer S-ACEs, handling the change of temporal length from 2 to 1 during training well. As for the reason why the model initialized as W/S/T=4/4/2 can perform better on accuracy very slightly. One explanation is that the initial temporal length does not affect the model performance while the target temporal length (or the actual temporal length) does, which is in line with the prior finding of [shen2024conventional] that the temporal length rarely affects the performance of multi-bit SNNs on static datasets. Therefore, for the saving of training time, we choose to set the initial temporal length as 1 in the experiments of the main text since the longer initial temporal length causes longer forward and backward inference time. Additionally, from another perspective, this could further prove that our multi-bit SNN has advanced its spatial processing ability to the limit, one time-step would suffice to fulfill the requirement of a high-accuracy inference.
In addition, we also naively apply our methods to Spikformer-8-512 on ImageNet with the same model hyperparameters as [shen2024conventional] and [zhou2022spikformer]. Their models achieve 70.87% [shen2024conventional] and 73.38% [zhou2022spikformer] with W/S/T = 4/2/1 and 16/1/4, respectively. While, our Spikformer achieves a higher 74.87% top-1 accuracy with lower bit budgets that W/S/T = 2.64/2.21/1.0.
@@
Secondly, as stated in Section A.5, we use the conventional sequential-processing pipeline to address the event dataset in the main text, i.e., the temporal squeezing and learnable $T$ are canceled. In this way, we can accurately control the temporal length and fully utilize the model-level temporal processing of SNN. This may raise the concern that the temporal squeezing could severely impair the temporal processing ability. As shown in Table A.2, we find the temporal squeezing can still outperform the quantization baseline [shen2024conventional] and maintain good performance under low time-steps, although the overall model performance, compared to the case without temporal squeezing, is decreased. One possible explanation is that although the model-level temporal processing ability may be diminished, the layer-wise temporal processing ability can still be maintained when using temporal squeezing. And, temporal squeezing also meets the temporal information dynamics that as the training goes on, more valuable and decisive information aggregates into the early time-steps [Kim_Li_Park_Venkatesha_Hambitzer_Panda_2023]. As listed in Table A.2, the best results of temporal squeezing show up under the low time-step settings. Temporal squeezing merely accelerates the aggregation process. That is why temporal squeezing would not severely undermine the sequential-processing ability and perform well even on the sequential dataset.
@@
Origin[shen2024conventional] Quant. Spf.[shen2024conventional] Our Spf. Our Spf. w/ Learnable $T_{l}$ W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 16/1/16 80.7 1/1/16 79.8 1.11/2.00/6.0 80.9 1.14/1.97/6.00 78.7 1/2/8 79.3 1.28/2.00/4.0 80.0 1.14/1.97/4.00 79.3 1/4/4 63.1 1.33/4.00/3.0 80.1 1.17/4.08/2.05 79.6 1/8/2 43.0 1.31/4.03/2.0 78.7 1.09/1.95/1.95 80.0 1/16/1 35.8 1.31/4.03/1.0 77.4 1.11/1.49/1.00 77.3
@@
A.8 Potential exploration to spiking audio task
We also conduct a simple experiment on the spiking audio dataset SHD [cramer2020heidelberg]. Specifically, we plug our method to the implementation of dilated convolution with learnable spacings (DCLS) [hammouamri2023learning], which is a synaptic-delay based framework to help SNNs operate on spiking audio tasks.
As shown in Table A.3, our method can help the original DCLS baseline to achieve better model accuracy, demonstrating a potential compatibility of our method with spiking audio task.
Additionally, we follow the exactly same hyperparameters and model architectures of DCLS [hammouamri2023learning] without changing a thing. The corresponding penalty coefficient $\lambda_{1}$ is set to 4e-2. The initial W is set to 6 and target Ws are set to 3 and 4 in Table A.3. The initial and target Ss are all set to 1. The initial and target Ts are all set to 140.
Table A.3: Extended experiments on the spiking audio dataset SHD.
Method Parameters (M) Top-1 DMUC [sun2024delayed] 0.24 91.48 CNN [cramer2020heidelberg] 1.01 92.40 RSNN [cramer2020heidelberg] 1.97 83.20 Attention RSNN [yao2021temporal] 0.14 91.08 RadLIF [bittar2022surrogate] 3.9 94.62 DL256-SNN-Dloss [sun2023learnable] 0.21 93.55 DCLS (T=140) [hammouamri2023learning] 0.21 95.07 Our DCLS-3.5/1/140 0.21 95.17 Our DCLS-4.4/1/140 0.21 95.54 "Our DCLS-x/x/x" denotes the learned bit widths of W/S/T.
A.9 Maintaining the asynchronous nature of SNNs
@@
Exploiting the spike perceptrons. Temporal squeezing itself is consistent with neurodynamics as it is essentially composed of a signal accumulation [gerstner2014neuronal] and a division operation. A similar accumulation mechanism can also be found in the current accumulation stage of spiking neuron, such as LIF and IF. Particularly, SpikeConverter neuron [liu2022spikeconverter] separates this accumulation from the spiking neuron modeling and uses a particular computing pipeline to handle this accumulation. Their accumulation is identical to ours, which is proven feasible in asynchronous computing. In addition, such accumulation has also been realized and implemented by neuromorphic circuits to process the DVS signals and generate the dynamic datasets [amir2017low, lin2021imagenet].
As such, we can simply use the same signal accumulation diagrams, i.e. the spike perceptrons, to realize temporal squeezing as what has been used in DVS cameras [li2017cifar10] and previous spiking neurons [liu2022spikeconverter]. In other words, such different computing paradigm as shown in Figure A.1 can be implemented under the existing neuromorphic computing and will not cause any loss of the asynchronous and clock-free nature.
@@
Using the two-stage pipeline or keeping single time-step. Although we have theoretically elaborated on the harmlessness of temporal squeezing to the asynchronous nature in the above, one may argue that the entire network’s spike flow could be interrupted by the squeezing stage, leading to longer inference latency. This concern has been discussed and resolved by [liu2022spikeconverter], using the two-stage pipeline. The first stage is spike accumulation, i.e., the temporal squeezing, the second stage is spike emission. These two separate stages can overlapped inter-layer. Specifically, assuming spike accumulation and emission both require 4 hardware clock cycles for computing. After the first cycle of the spike emission of the previous layer, the first computing cycle of the spike accumulation of the next layer can be started instead of waiting for the whole 4 cycles of the spike emission of the previous layer being finished. Thus, 3 cycles are overlapped, and the actual latency of the two-stage spiking pipeline is only 5 cycles, only one more cycle that the ordinary spiking pipeline. Therefore, using temporal squeezing only introduces very limited latency increase.
A.10 Overall algorithm of the proposed bit allocation